A small 2.00 g plastic ball is suspended by a 20.0 cm long string in a uniform electric field. If the ball is in equilibrium when the string makes a 15.0 degree angle with the vertical, what is the net charge on the ball?
1
Expert's answer
2020-10-20T07:14:44-0400


From the picture we can write that

"T\cos 15^o=mg" ......(1)

"T\sin15^o=qE" ......(2)


From 1 and 2

"\tan 15^o=\frac{qE}{mg}"

"q=0.27\times0.002\times10/E=0.0054/E"